Researchers enrolled 2,100 women in a study and followed them every year for four years to determine the incidence rate of heart disease. After the first year, none had a new diagnosis of heart disease, but 150 were lost to follow-up. After year two, three had a new diagnosis of heart disease, and another 122 had been lost to follow-up. After year three, another eighteen had new diagnoses of heart disease, and 602 had been lost to follow-up. After four years, an additional 15 had new diagnoses with heart disease, and 258 more had been lost to follow-up. Calculate the incidence rate of heart disease among this cohort. Assume that persons with new diagnoses of heart disease and those lost to follow-up were disease-free for half the year, and thus contribute ½ year to the denominator. (Show your work and label your calculations)
Answer #1
Year Under risk new diagnosis lost numerator denominator Incidence rate
1 2100 0 150 0 2100-150 +150/2=2025 0
2 2100-150 =1950 3 122 3 1950-122+122/2 +3/2=1890.5 0.0016
3 1950-122-3=1825 18 602 18 1825-602+602/2 +18/2= 1533 0.0117
4 1825-602-18= 1205 15 258 15 1205-258+258/2+15/2 = 1083.5 0.0138

Note : In year 1 , no new diagnosis , thus numerator =0

For denominator , we subtract num of lost cases from population under study and then add half of the lost cases =2100-150+150/2 = 2025

In year 2 , three new diagnosis , thus numerator =3

For denominator , we subtract num of lost cases from remaining population under study and then add half of the lost cases and half of the cases with new diagnosis =1950-122+122/2 +3/2=1890.5

and so on